OpenStudy (anonymous):

solve by factoring. 2x^2-3x+1=0

jimthompson5910 (jim_thompson5910):

2x^2-3x+1=0 (2x-1)(x-1) = 0 2x-1= 0 or x-1=0 I'll let you finish up

OpenStudy (anonymous):

i cant finish up i dont understand factoring i tried i just dont get it.

jimthompson5910 (jim_thompson5910):

from here, you solve each for x 2x-1= 0 or x-1=0 2x= 1 or x=1 x = 1/2 or x = 1

jimthompson5910 (jim_thompson5910):

So the two solutions are x = 1/2 or x = 1

OpenStudy (anonymous):

thank youu (:

jimthompson5910 (jim_thompson5910):

yw